OpenStudy (anonymous):

Use the discriminant to determine how many x-intercepts the graph of the equation has. y = – 4x2 + 3x + 2 A. zero B. one C. two D. three

OpenStudy (anonymous):

do you know the formula of discriminant?

OpenStudy (anonymous):

the discriminant is 41

OpenStudy (anonymous):

not really, had to look that up.

OpenStudy (anonymous):

no it's okay you don't have to look that's what I'm here for lol

OpenStudy (anonymous):

alright \[D = b^2 - 4ac \]

OpenStudy (anonymous):

that is the formula

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what you've been give is a quadratic formula you know how a quadratic formula looks like right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

alrighty if you click on this link, you will know what exactly are those a, b and c are in the Discriminant formula http://www.gradeamathhelp.com/image-files/quadratic-formula-abc.gif

OpenStudy (anonymous):

either my brain aint working or that doesn't look right?

OpenStudy (anonymous):

probably me

OpenStudy (anonymous):

mmm which part is it that you're having difficulty?

OpenStudy (anonymous):

do you mean a, b, c are in discriminent form?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

a,b and the c are the coefficients

OpenStudy (anonymous):

coefficients meaning any number that is in front of the x

OpenStudy (anonymous):

like 2 is the coefficient in 2x

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so now look at the coefficients in your equation and tell me what are they?

OpenStudy (anonymous):

\[y = -4x^2 + 3x + 2\]

OpenStudy (anonymous):

4, 2, & 3?

OpenStudy (anonymous):

yupppp! so which one is the a and the b and the c? (according to the link I gave you)

OpenStudy (anonymous):

4 = a, 3 = b, and 2 = c?

OpenStudy (anonymous):

yup!! you're doing very good :)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

now if we plug in the numbers into our discriminant formula which is \[D = b^2 - 4ac\]

OpenStudy (anonymous):

we will get \[D = 3^2 - (-4*4*2)\]

OpenStudy (anonymous):

D equals???

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so what do we get for D?

OpenStudy (anonymous):

41

OpenStudy (anonymous):

Correct!!! :D

OpenStudy (anonymous):

now to answer your main question a quadratic equation will always give you a graph that looks like U (the way how you know it's a quadratic if it has a \[x^2 \] and a x

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

and if your D is more than zero there are two places that graph intersect like this: